OpenStudy (anonymous):

If r(t)=−4ti+4t2j+2tk, compute the tangential and normal components of the acceleration vector.

OpenStudy (amistre64):

i do need the practice

OpenStudy (anonymous):

Well practice on. If you get it by 9:45 I can tell you if you are correct

OpenStudy (amistre64):

lol....

OpenStudy (amistre64):

have to be able to get the site to run to right on my computer tho

OpenStudy (amistre64):

\[ r(t)=<−4t,4t^2,2t>\] \[ r'(t)=<−4,8t,2>\] \[ |r'(t)|=\sqrt{16+64t^2+4}\] \[ r''(t)=<0,8,0>\] \[ r'.r''=<−4,8t,2>.<0,8,0>=64t\] \[ r'(x)r''=<−4,8t,2>(x)<0,8,0>=<-16,0,-32>\] \[ |r'(x)r''|=\sqrt{16^2+32^2}\] \[a_t=\frac{r'.r''}{|r'|}=\frac{64t}{\sqrt{20+64t^2}}\] \[a_n=\frac{|r'(x)r''|}{|r'|}=\frac{\sqrt{16^2+32^2}}{\sqrt{20+64t^2}}\]